# Scaling Cross-Embodiment World Models for Dexterous Manipulation

### v1

@@ I Introduction
Cross-embodiment learning underpins the vision of building generalist embodied agents that operate across diverse physical embodiments by learning from heterogeneous data [3, 46]. This paradigm is becoming increasingly important as variations in morphology arise naturally from manufacturing tolerances, wear, and hardware upgrades, while large-scale robot deployments generate diverse cross-embodiment datasets (e.g., [22, 11]). Prior progress has shown embodiment-level generalization in locomotion [3] and in manipulation with parallel grippers [4, 17, 43, 23], whereas dexterous manipulation has largely been limited to grasping [42, 10] and in-hand reorientation [25]. Extending such generalization to broader task domains, such as non-prehensile object pushing and deformable object manipulation, remains challenging due to complex object dynamics and the need for fine-grained contact control.
@@
Dexterous hands offer a case for cross-embodiment learning. Despite the challenge of contact-rich, high-DoF control, multifingered robot hands are morphologically similar to one another and to human hands. This anthropomorphism suggests that cross-embodiment datasets, including not only robot datasets but also human–object interactions, can be mutually informative. This raises two central questions: what invariant knowledge about the external world and contact interactions underlies purposeful action across distinct kinematics and control spaces, and how should state and action be represented so that human and robot data become jointly useful? We propose that world models, learned dynamics models that predict future states given the current state and action [1, 37], provide such an interface by capturing environment dynamics invariant to embodiment. The central challenge then becomes designing state and action representations that are embodiment-agnostic, mapping heterogeneous data and control spaces into a shared domain where a single predictive model can learn from all embodiments and enable model-based control across diverse kinematic structures.
@@
To this end, we represent both human and robot hands as particles (i.e., 3D point sets), with actions defined as particle displacements. A graph-based dynamics model [33, 2, 34, 35, 51] predicts particle motion while exploiting spatial locality and equivariance. We co-train the model on simulated robot–object interaction data and real human–object interaction data and study how generalization of the learned model scales with the morphological diversity in the training domain. For control, we sample robot joint actions, which are then converted to particle action representation via forward kinematics, to enable model-predictive control in the particle space. This action abstraction unifies control problems across embodiments, allowing the learned model to be deployed on hands with varied control spaces without motion retargeting or expert demonstration collection.
@@
We evaluate learned world models both in simulation and on real hardware. In simulation, we observe an embodiment scaling trend [3]: training on more simulated hands consistently improves generalization to unseen embodiments. In the real world, we find that models trained solely on human data can already transfer to robotic hands despite the embodiment gap, and that incorporating an appropriate amount of simulation data further improves both predictive accuracy and planning performance. Our best model, co-trained on simulated robot data and real human data, enables a 6-DoF PSYONIC Ability Hand and a 12-DoF Robot Era XHand to successfully perform deformable object manipulation. These results demonstrate the promise of world models as a unifying interface for cross-embodiment learning and generalization in robot manipulation.
@@ II-A Cross-Embodiment Learning
One key goal of cross-embodiment learning is to learn from diverse embodiments, with learning from human data being a special case. Most existing approaches are model-free, which learn mappings from observations to actions end-to-end via reinforcement learning (RL) or behavior cloning (BC). Prior work includes using human demonstrations to guide RL for dexterous manipulation [27, 6], cross-embodiment RL training for locomotion [5], and scaling to many embodiments through a combination of RL and BC [3, 25]. When high-quality demonstrations are available, BC can provide stable learning signals, which powered recent efforts on training vision-language-action (VLA) models on cross-embodiment datasets [4, 17, 23, 54, 49]. Human demonstrations can also be adapted to robots via motion retargeting for anthropomorphic robots [28]. These works either require extensive RL training in simulation, limited to domains where the sim-to-real gap is moderate, or require expert demonstrations in the real world.
Model-based approaches instead learn world models that explicitly predict the effects of actions from interaction data, and have shown strong performance in tasks requiring fine-grained control [1]. One related study explored particle-based models from human–object interactions [14], but the resulting policy was too inaccurate for direct model-based control, and transfer to novel embodiments remained infeasible. In this work, we aim to learn accurate cross-embodiment world models through embodiment scaling, using data from both robotic and human hands. We show that, while using human-only data already achieves moderate performance, incorporating it with diverse simulation cross-embodiment data further enhances performance.
@@ II-B World Model Learning for Robotic Manipulation
World models, predictive models that forecast the effects of actions, are central to model-based robotic control [1]. Their effectiveness depends heavily on state representation. Pixel-based models [9, 48, 8] can exploit large-scale visual data but struggle to produce physically accurate predictions in contact-rich settings with extensive training. Particle-based models are more physically grounded and incorporate stronger inductive biases. They have enabled manipulation of clothes [20, 38], ropes [52], granular object piles [41], and plasticine [34, 33], but are often trained on single-embodiment data. This work extends particle-based world model learning to multi-embodiment training, including both robotic and human hands, and systematically studies its transferability to unseen embodiments.
@@ II-C Dexterous Manipulation
Dexterous manipulation is a long-standing challenge in robotics [30, 21], largely due to the high degrees of freedom of multi-fingered hands and complex contact patterns. Classic control methods rely on analytical object models [29, 36, 24], which may not capture hard-to-model factors, such as frictional contact or actuator drift, and can be hard to obtain for deformable objects. Recent learning-based approaches, including RL and BC, have shown success in rigid object manipulation, such as grasping [40, 10], in-hand reorientation [13, 26], and tool use [32, 50, 39]. However, manipulating deformable objects using multi-fingered hands remains under-explored [18, 53], due to the combined challenges of high-dimensional control and complex object dynamics. In this work, we learn a model of environment dynamics and integrate it with model-predictive control. By introducing embodiment-agnostic state and action representations, we enable learning from both robotic and human data, allowing a single model to operate over diverse dexterous hands.
@@ III-B Perception Module
The perception module performs state estimation for data collection and deployment. We use a multi-view camera setup [2, 33, 34, 38]. Cameras are placed at fixed positions around the scene so that each captures the interaction from a different viewpoint, ensuring comprehensive coverage.
For real-world human data collection, we reconstruct hand meshes from the multi-view images using POEM-v2 [47], and sample particles with farthest point sampling (FPS). For object perception, we fuse multi-view point clouds, perform Poisson surface reconstruction to obtain a smooth surface [16], and apply FPS. In both cases, FPS allows us to obtain particles that preserve the full geometry of the meshes.
@@ III-C World Model Architecture
We consider adopting graph neural networks (GNNs) as our world model architecture, as the locality and equivariance are useful inductive biases [15] that allow the learned model to generalize to objects and hands with different shapes. We use DPI-Net [19], a GNN that models local particle interactions through message passing and captures global effects via multi-step hierarchical propagation.
@@ III-D Model-Based Planning
Inspired by the insight that human hand motions lie in low-dimensional manifolds of the full configuration space [12], we design action spaces for efficient planning, and use the cross-entropy method for model-predictive control (MPC).
@@ IV-A Experimental Setup
Task setup. We consider two representative dexterous manipulation tasks: non-prehensile rigid object pushing [2, 7] and deformable object reshaping [2, 33, 34]. In Object Pushing, the goal is to reorient a box to a target orientation. In Plasticine Reshaping, the goal is to mold plasticine into a target shape specified by a point cloud. Both tasks require precise contact control and reasoning about object dynamics.
Simulation setup. We simulate six dexterous hands representative of commonly used multi-fingered designs: Ability Hand (6-DoF), Allegro Hand (16-DoF), XHand (12-DoF), Leap Hand (16-DoF) [31], Shadow Hand (24-DoF), and a URDF variant of the Shadow Hand without its forearm (24-DoF). For the rigid-body task (Object Pushing), we use SAPIEN [44] for data collection. For deformable object manipulation, we use the Rewarped simulation platform [45], a differentiable multiphysics simulator. We collect 100 trajectories per task, where the robots perform random actions in the predefined action space.
@@ IV-B Evaluating Cross-Embodiment World Model Learning
  • Embodiment scaling law [3]: Prediction error decreases as more embodiments are included, and variance across subsets shrinks, indicating more stable models with broader embodiment diversity.

### v2

@@ I Introduction
Cross-embodiment learning underpins the vision of building generalist embodied agents that operate across diverse physical embodiments by learning from heterogeneous data [ai2025towards, Yang2024pushing]. This paradigm is becoming increasingly important as variations in morphology arise naturally from manufacturing tolerances, wear, and hardware upgrades, while large-scale robot deployments generate diverse cross-embodiment datasets (e.g., [openx, rh20t]). Prior progress has shown embodiment-level generalization in locomotion [ai2025towards] and in manipulation with parallel grippers [Black2024pizero, Kim2024openvla, cage, octo_2023], whereas dexterous manipulation has largely been limited to grasping [dro, fang2025anydexgrasp] and in-hand reorientation [getzero]. Extending such generalization to broader task domains, such as non-prehensile object pushing and deformable object manipulation, remains challenging due to complex object dynamics and the need for fine-grained contact control.
@@
Dexterous hands offer a case for cross-embodiment learning. Despite the challenge of contact-rich, high-DoF control, multifingered robot hands are morphologically similar to one another and to human hands. This anthropomorphism suggests that cross-embodiment datasets, including not only robot datasets but also human–object interactions, can be mutually informative. This raises two central questions: what invariant knowledge about the external world and contact interactions underlies purposeful action across distinct kinematics and control spaces, and how should state and action be represented so that human and robot data become jointly useful? We propose that world models, learned dynamics models that predict future states given the current state and action [ai2025review, tenenbaum2011howto], provide such an interface by capturing environment dynamics invariant to embodiment. The central challenge then becomes designing state and action representations that are embodiment-agnostic, mapping heterogeneous data and control spaces into a shared domain where a single predictive model can learn from all embodiments and enable model-based control across diverse kinematic structures.
@@
To this end, we represent both human and robot hands as particles (i.e., 3D point sets), with actions defined as particle displacements. A graph-based dynamics model [robocook, robopack, robocraft, shi2023robocraft, zhang2024adaptigraph] predicts particle motion while exploiting spatial locality and equivariance. We co-train the model on simulated robot–object interaction data and real human–object interaction data and study how generalization of the learned model scales with the morphological diversity in the training domain. For control, we sample robot joint actions, which are then converted to particle action representation via forward kinematics, to enable model-predictive control in the particle space. This action abstraction unifies control problems across embodiments, allowing the learned model to be deployed on hands with varied control spaces without motion retargeting or expert demonstration collection.
@@
We evaluate learned world models both in simulation and on real hardware. In simulation, we observe an embodiment scaling trend [ai2025towards]: training on more simulated hands consistently improves generalization to unseen embodiments. In the real world, we find that models trained solely on human data can already transfer to robotic hands despite the embodiment gap, and that incorporating an appropriate amount of simulation data further improves both predictive accuracy and planning performance. Our best model, co-trained on simulated robot data and real human data, enables a 6-DoF PSYONIC Ability Hand and a 12-DoF Robot Era XHand to successfully perform deformable object manipulation. These results demonstrate the promise of world models as a unifying interface for cross-embodiment learning and generalization in robot manipulation.
@@ II-A Cross-Embodiment Learning
One key goal of cross-embodiment learning is to learn from diverse embodiments, with learning from human data being a special case. Most existing approaches are model-free, which learn mappings from observations to actions end-to-end via reinforcement learning (RL) or behavior cloning (BC). Prior work includes using human demonstrations to guide RL for dexterous manipulation [dexmv, chen2024vividex], cross-embodiment RL training for locomotion [bohlinger2024onepolicy], and scaling to many embodiments through a combination of RL and BC [ai2025towards, getzero]. When high-quality demonstrations are available, BC can provide stable learning signals, which powered recent efforts on training vision-language-action (VLA) models on cross-embodiment datasets [Black2024pizero, Kim2024openvla, octo_2023, rt-2, egovla]. Human demonstrations can also be adapted to robots via motion retargeting for anthropomorphic robots [qiu2025-humanpolicy]. These works either require extensive RL training in simulation, limited to domains where the sim-to-real gap is moderate, or require expert demonstrations in the real world.
Model-based approaches instead learn world models that explicitly predict the effects of actions from interaction data, and have shown strong performance in tasks requiring fine-grained control [ai2025review]. One related study explored particle-based models from human–object interactions [hong2025learning], but the resulting policy was too inaccurate for direct model-based control, and transfer to novel embodiments remained infeasible. In this work, we aim to learn accurate cross-embodiment world models through embodiment scaling, using data from both robotic and human hands. We show that, while using human-only data already achieves moderate performance, incorporating it with diverse simulation cross-embodiment data further enhances performance.
@@ II-B World Model Learning for Robotic Manipulation
World models, predictive models that forecast the effects of actions, are central to model-based robotic control [ai2025review]. Their effectiveness depends heavily on state representation. Pixel-based models [ebert2018visual, yang2023learning, du2023learning] can exploit large-scale visual data but struggle to produce physically accurate predictions in contact-rich settings with extensive training. Particle-based models are more physically grounded and incorporate stronger inductive biases. They have enabled manipulation of clothes [Lin2021learningvisible, tian2025uniclothdiff], ropes [zhang2024dynamics], granular object piles [wang2023dynamic], and plasticine [robocraft, robocook], but are often trained on single-embodiment data. This work extends particle-based world model learning to multi-embodiment training, including both robotic and human hands, and systematically studies its transferability to unseen embodiments.
@@ II-C Dexterous Manipulation
Dexterous manipulation is a long-standing challenge in robotics [salisbury1982articulated, mordatch2012contact], largely due to the high degrees of freedom of multi-fingered hands and complex contact patterns. Classic control methods rely on analytical object models [rus1999hand, suh2025dexterous, Pang2023global], which may not capture hard-to-model factors, such as frictional contact or actuator drift, and can be hard to obtain for deformable objects. Recent learning-based approaches, including RL and BC, have shown success in rigid object manipulation, such as grasping [wan2023unidexgrasp++, fang2025anydexgrasp], in-hand reorientation [handa2023dextreme, qi2023hand], and tool use [shaw2024bimanual, yin2025dexteritygen, wan2025lodestar]. However, manipulating deformable objects using multi-fingered hands remains under-explored [dexdeform, zhaole2024dexdlo], due to the combined challenges of high-dimensional control and complex object dynamics. In this work, we learn a model of environment dynamics and integrate it with model-predictive control. By introducing embodiment-agnostic state and action representations, we enable learning from both robotic and human data, allowing a single model to operate over diverse dexterous hands.
@@ III-B Perception Module
The perception module performs state estimation for data collection and deployment. We use a multi-view camera setup [robopack, robocook, robocraft, tian2025uniclothdiff]. Cameras are placed at fixed positions around the scene so that each captures the interaction from a different viewpoint, ensuring comprehensive coverage.
For real-world human data collection, we reconstruct hand meshes from the multi-view images using POEM-v2 [poem-v2], and sample particles with farthest point sampling (FPS). For object perception, we fuse multi-view point clouds, perform Poisson surface reconstruction to obtain a smooth surface [kazhdan2006poisson], and apply FPS. In both cases, FPS allows us to obtain particles that preserve the full geometry of the meshes.
@@ III-C World Model Architecture
We consider adopting graph neural networks (GNNs) as our world model architecture, as the locality and equivariance are useful inductive biases [Leslie2020thefoundation] that allow the learned model to generalize to objects and hands with different shapes. We use DPI-Net [dpi-net], a GNN that models local particle interactions through message passing and captures global effects via multi-step hierarchical propagation.
@@ III-D Model-Based Planning
Inspired by the insight that human hand motions lie in low-dimensional manifolds of the full configuration space [grasp_taxonomy], we design action spaces for efficient planning, and use the cross-entropy method for model-predictive control (MPC).
@@ IV-A Experimental Setup
Task setup. We consider two representative dexterous manipulation tasks: non-prehensile rigid object pushing [robopack, chi2024diffusion] and deformable object reshaping [robopack, robocook, robocraft]. In Object Pushing, the goal is to reorient a box to a target orientation. In Plasticine Reshaping, the goal is to mold plasticine into a target shape specified by a point cloud. Both tasks require precise contact control and reasoning about object dynamics.
Simulation setup. We simulate six dexterous hands representative of commonly used multi-fingered designs: Ability Hand (6-DoF), Allegro Hand (16-DoF), XHand (12-DoF), Leap Hand (16-DoF) [leaphand], Shadow Hand (24-DoF), and a URDF variant of the Shadow Hand without its forearm (24-DoF). For the rigid-body task (Object Pushing), we use SAPIEN [Xiang_2020_SAPIEN] for data collection. For deformable object manipulation, we use the Rewarped simulation platform [rewarped], a differentiable multiphysics simulator. We collect 100 trajectories per task, where the robots perform random actions in the predefined action space.
@@ IV-B Evaluating Cross-Embodiment World Model Learning
  • Embodiment scaling law [ai2025towards]: Prediction error decreases as more embodiments are included, and variance across subsets shrinks, indicating more stable models with broader embodiment diversity.